OpenStudy (anonymous):

Use the slope formula to find the missing coordinate. P(3, 2), Q(8,_ ) m =

OpenStudy (anonymous):

Hi savanna, so do you know the slope formula?

OpenStudy (amistre64):

i think theres a value for "m" thats missing

OpenStudy (anonymous):

yeah its 3/5

OpenStudy (amistre64):

subtract your points and see how your values would math the given slope parts is my thought

OpenStudy (amistre64):

P(3,2) -Q(8,n) ------- 5, 2-n; slope = \(\frac{2-n}{5}=\frac{3}{5}\)

OpenStudy (amistre64):

that should be a -5, my bad ;)

OpenStudy (anonymous):

thanks i get it now :)

OpenStudy (amistre64):

good job ;)

OpenStudy (anonymous):

Awesomely done amistre ;)